OpenStudy (anonymous):

Given the function g(x) = −2x + 4, find g(−5). −6 4.5 6 14

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

it not

OpenStudy (gg):

g(-5)=-2(-5)+4=10+4=14

OpenStudy (anonymous):

yeah, it is, unless you gave me the wrong function?

OpenStudy (anonymous):

oooops...I misread it. I gave you g'(-5) not g(-5) lol

OpenStudy (anonymous):

i dint

OpenStudy (anonymous):

yeah 14 is correct for g(-5)